OpenStudy (anonymous):

if trying to find all solutions on the interval 0,2pi sin2x=sinx so x=1/2? or how do i solve?

myininaya (myininaya):

sin2x=2sinxcosx so 2sinxcosx=sinx 2sinxcosx-sinx=0 sinx(2cosx-1)=0 sinx=0 and 2cosx-1=0

myininaya (myininaya):

lets solve sinx=0 first

myininaya (myininaya):

what x makes sinx zero

myininaya (myininaya):

think about your unit circle

myininaya (myininaya):

sin(0) is what?

OpenStudy (anonymous):

0 or pi

myininaya (myininaya):

sin(0)=0 sin(2pi)=0 so x=0 and 2pi now we have the other equation 2cosx-1=0 => cosx=1/2

myininaya (myininaya):

what can plug in where the x is that give you 1/2 (we are looking at the equation cosx=1/2)

OpenStudy (anonymous):

60 degrees

OpenStudy (anonymous):

the other btw was 0 degrees or 180 degrees for sin

myininaya (myininaya):

oh yes and pi very good for the first

OpenStudy (anonymous):

so answeres are 0,1/2,pi?

myininaya (myininaya):

so 60 degrees and what else for this equation equation

myininaya (myininaya):

no not 1/2

myininaya (myininaya):

60 degrees or pi/3 there is also another one for the second equation

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

0,pi/3,pi that seems like all we can get i think cause thats both sin and hte one cos

OpenStudy (anonymous):

@myininaya: Congrats for the new title. I was you 100th fan by the way :D

OpenStudy (anonymous):

i was 99th XD

myininaya (myininaya):

thanks lol

myininaya (myininaya):

so anyways what is cos(300 degrees)

OpenStudy (anonymous):

10pi/6

myininaya (myininaya):

no maybe it will help if i wrote it in radians what is cos(5pi/3)

OpenStudy (anonymous):

isnt that the same just reduced?

myininaya (myininaya):

ok 10pi/6=5pi/3 but what is cos(5pi/3)

myininaya (myininaya):

=1/2

OpenStudy (anonymous):

-1/2?

OpenStudy (anonymous):

isnt it - cause bottom right quadrent?

OpenStudy (anonymous):

nvm cos is x so positive if on right side

myininaya (myininaya):

k :) so the solutions to sinx=0 in the interval [0,2pi] are 0,pi,2pi and the solutions to cosx=1/2 in the interval [0,2pi] are pi/3, 5pi/3

myininaya (myininaya):

you have 5 solutions in all

OpenStudy (anonymous):

thank you

myininaya (myininaya):

np